PROBLEM: A key step in the production of sulfuric acid is the oxidation of S*O_{2}(g) to S*O_{3}(g) :



2S*O_{2}(g) + O_{2}(g) -> 2S*O_{3}(g)



At 298 K, Delta*G = - 141.6kJ; Delta*H = - 198.4kJ and Delta*S = - 187.9J / K



(a) Use the data to decide if this reaction is spontaneous at 25 degrees * C , and predict how Delta*G will change with increasing T.



(b) Assuming Delta*H and Delta*S are constant with increasing T, is the reaction spontaneous at 900.^ C ?

1
Expert's answer
2022-06-01T07:50:03-0400

2SO2 + O2 = 2SO3

∆G = -141,6 kJ

∆H = -198,4 kJ

∆S = -187,9 J = -0,1879 kJ

a)

Since ∆G is negative at 25 ° C (298K), the reaction occurs spontaneously.

According to the formula ∆G = ∆H-T*∆S , an increase in temperature changes ∆G in a positive direction. therefore, it does not occur spontaneously. or the reaction goes in the opposite direction.

b)

∆H = -198,4 kJ

∆S = -0,1879 kJ

T= 900°C = 1173 K

∆G = ∆H - T∆S = -198,4 - (1173*(-0,1879))= 22,0067 ≈ 22 kJ

Since ∆G> 0 is present at 900 ° C, the reaction does not proceed spontaneously.